# Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity

### v1

@@ Abstract
We investigate the transition rates of a centripetally accelerated atom interacting with electromagnetic vacuum fluctuations inside a high-quality cavity. Our findings reveal that the emission and excitation rates can be extensively manipulated by adjusting the cavity’s normal mode frequency and the rotational angular velocity. Using experimentally feasible parameters, we demonstrate that, in one scenario, the excitation rate can reach magnitudes as high as $10^{7}~{}{\rm s}^{-1}$, while the emission rate remains negligible, trailing by 9 orders of magnitude. This suggests the potential for substantial population inversion in an ensemble of atoms. In another scenario, both the emission and excitation rates can simultaneously reach magnitudes as high as $10^{7}~{}{\rm s}^{-1}$, indicating that millions of transitions per second are expected even for a single atom. These remarkable results not only offer a new method for controlling the radiative properties of atoms but also open avenues for the experimental verification of the circular Unruh effect using cutting-edge quantum technologies.
@@
Introduction.— The study of atomic transitions has long been central to our understanding of the interaction between matter and electromagnetic fields within the framework of quantum electrodynamics. In free space, an atom in its excited state can spontaneously emit a photon and transition to a lower energy state, a process known as spontaneous emission. Conversely, spontaneous excitation, where an atom in its ground state absorbs energy from the surrounding electromagnetic vacuum to transition to an excited state, is not observed for inertial atoms in free space. It is well understood that spontaneous emission in vacuum is governed by the interaction between the atom and the fluctuating electromagnetic field. Therefore, if vacuum fluctuations are somehow modified, changes in the atom’s radiative properties would naturally occur.
One particularly intriguing scenario arises when atoms are placed in a resonant cavity, as the presence of the cavity alters the density of states available for the fluctuating electromagnetic field. This modification can lead to significant changes in the transition rates of the atom, exemplified by the Purcell effect Purcell46 , first proposed in 1946, where the rate of spontaneous emission is enhanced due to the modified vacuum environment within the cavity. This enhancement was later observed in both the microwave Purcell2 and optical regimes Purcell3 . The Purcell effect has been instrumental in developing a deeper understanding of how confined fluctuating electromagnetic fields in vacuum can influence atomic behavior, opening up new avenues for manipulating and controlling quantum systems and light-matter interactions.
@@
Vacuum fluctuations of the electromagnetic field can be profoundly altered not only by confinement to bounded spaces, such as cavities, but also by the non-inertial motion of atoms. Atoms in such motion interact with these modified vacuum fluctuations in unique ways, leading to phenomena that have no counterpart in inertial motion. For instance, an atom undergoing uniform linear acceleration perceives the vacuum, as seen by an inertial observer, as a thermal bath. This leads to the atom spontaneously transitioning to a higher energy level, a phenomenon known as the Unruh effect Davies75 ; Unruh76 ; Fulling73 . Additionally, the emission rate of an accelerated atom is also enhanced compared to that of an inertial atom, mimicking the behavior of an atom in a thermal environment. However, the excitation rate typically remains much lower than the emission rate, except under extremely high accelerations where it can approach the emission rate. Therefore, while the Unruh effect is theoretically compelling, it has remained elusive in direct experimental observation due to the extreme accelerations required.
@@
Similar to the case of linear acceleration, an observer undergoing uniform centripetal acceleration would also perceive the vacuum of an inertial atom as an environment with radiation, although the spectrum of this radiation is non-thermal Bell83 ; Hacyan1986 ; Bell87 ; Kim1987 ; Unruh98 ; Rosu2005 . Consequently, spontaneous excitation can occur for atoms under centripetal acceleration in a vacuum, a phenomenon that can be regarded as the circular Unruh effect. The emission rate in this scenario can also be enhanced by centripetal acceleration, although the behavior differs significantly from that under uniform linear acceleration Letaw1981 ; Takagi1984 ; Letaw1980 ; Rogers88 ; Davies1996 ; Lorenci2000 ; YaoJin2014-1 ; YaoJin2014-2 ; Biermann2020 ; Bunney23 ; Bunney23-2 ; yu24 .
@@
Interestingly, recent studies have shown that when the rotational angular velocity exceeds the transition frequency of the atom, the excitation rate can reach the same order of magnitude as the emission rate, even when the orbital radius and thus the centripetal acceleration are extremely small yu24 . These findings not only challenge the conventional understanding that a significant Unruh effect requires a large acceleration and reveal that the radiative properties of a rotating atom can be profoundly different from those of inertial atoms, even at vanishingly small centripetal accelerations, but also naturally lead to further questions. Specifically, on one hand, how would the radiative properties of inertial atoms inside a cavity change if they are rotated? Can the maximum cavity-amplified emission rate for inertial atoms be further enhanced by centripetal acceleration, as it is in free space? On the other hand, how would the radiative properties of rotating atoms in free space change if they are placed inside a cavity? Can the excitation rate of non-inertial atoms be drastically amplified by cavities, similar to how the emission rate is for inertial atoms? If possible, can the excitation rate even exceed the emission rate, leading to population inversion? These questions are central to the present study, where we explore the interplay between the effects of atomic circular acceleration, a form of non-inertial motion, and the confinement of the electromagnetic field within the cavity, with an eye toward controlling atomic radiative properties and experimentally verifying the circular Unruh effect using cutting-edge quantum technologies.
@@
Transition rates of centripetally accelerated atoms in a cavity.— We consider a two-level atom with a proper transition frequency $\omega_{0}$ coupled to fluctuating vacuum electromagnetic fields within a cavity. Assuming that the atom moves along a circular orbit, its trajectory in the laboratory frame is described by
$x(t)=R\cos(\Omega t),~{}y(t)=R\sin(\Omega t),~{}z(t)=0\;,$ (1)
where $R$ represents the radius of the orbit, $\Omega$ the rotational angular velocity, and $t$ the coordinate time. In the laboratory frame, the Hamiltonian characterizing the dipole interaction between the atom and the electromagnetic field can be expressed as
$H_{I}=-\sum_{m=\rho,\phi,z}D_{m}\mathcal{E}_{m}\;,$ (2)
where $D_{m}$ is the electric dipole moment operator of the atom, and the components of $\mathcal{E}_{m}$ are
@@
Using perturbation theory, the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$ of the atom in the laboratory frame can be derived as follows
[…]
where $t_{-}=t-t^{\prime}$, ${d}_{m}=\langle e|D_{m}|g\rangle$ is the dipole transition matrix element with $|g\rangle$ ($|e\rangle$) being the ground (excited) state of the atom, $G_{mn}(t_{-})=\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{n}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$ are the two-point functions of the electromagnetic fields with $|0\rangle$ being the vacuum state of the electromagnetic fields, and $\omega=\omega_{0}(1-{\Omega^{2}R^{2}}/{c^{2}})^{\frac{1}{2}}$ is the transition frequency of the atom in the laboratory frame. The $+$ and -- signs in the exponential function correspond to the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$, respectively. For simplicity, if we assume the transition matrix elements ${d}_{m}$ are real, then the cross terms in Eq. (4) vanish. For the explicit forms of the two-point functions $\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{m}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$. These functions are dependent on the explicit form of the density of states within the cavity $\rho(\omega_{k})$, where $\omega_{k}$ denotes the frequency of the field mode. The density of states within a cavity is typically modeled as a Lorentzian distribution given by
$\rho(\omega_{k})=\frac{1}{\pi}\frac{\omega_{c}/Q}{(\omega_{c}/Q)^{2}+(\omega_{k}-\omega_{c})^{2}}\;,$ (5)
where $Q~{}(Q\gg 1)$ represents the quality factor, and $\omega_{c}$ denotes the normal mode frequency of the cavity. By substituting the explicit forms of the field correlation functions into Eq. (4), the leading terms of the emission rate $\Gamma_{\downarrow}$ in the non-relativistic limit, i.e., when the linear velocity $v=R\Omega$ is much smaller than the speed of light $c$, can be calculated as
@@
$+\left(d_{\rho}^{2}+d_{\phi}^{2}\right)\times\left[\frac{\omega_{0}+\Omega}{Q^{2}\left(\omega_{0}+\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right.$
$\left.+\frac{\left(\omega_{0}-\Omega\right)\Theta\left(\omega_{0}-\Omega\right)}{Q^{2}\left(\omega_{0}-\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right]\bigg{\}}.$ (6)
@@
$\Gamma_{\uparrow}\approx\frac{Q\omega_{c}\left(\Omega-\omega_{0}\right)\left(d_{\rho}^{2}+d_{\phi}^{2}\right)}{6\epsilon_{0}\hbar V\left[Q^{2}\left(\Omega-\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}\right]},$ (7)
for $\Omega>\omega_{0}$, and
$\Gamma_{\uparrow}\approx\frac{QR^{2}\omega_{c}\left(2\Omega-\omega_{0}\right){}^{3}\left(d_{\rho}^{2}+d_{\phi}^{2}\right)}{40c^{2}\epsilon_{0}\hbar V\left[Q^{2}\left(2\Omega-\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}\right]},$ (8)
for $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$. Here $\epsilon_{0}$ is the vacuum permittivity, $V$ represents the volume of the cavity, $\Theta(x)$ is the Heaviside step function, which is equal to 00 when $x<0$ and equal to $1$ when $x>0$. Note that here we do not consider the case when $0<\Omega\leq\frac{\omega_{0}}{2}$, as the leading term of the excitation rate is of order $({R\Omega}/{c})^{4}$ or higher. In the following, we discuss explicitly the properties of the transition rates for centripetally accelerated atoms within a cavity, focusing on the interplay between the effects of centripetal acceleration and cavity. It should be noted that the analysis assumes $Q\gg 1$.
Rotational effect.— First, we examine the impact of rotation on the radiative properties of atoms by comparing the emission rate of centripetally accelerated atoms inside a cavity with that of inertial atoms in the same environment. In particular, we focus on whether the maximum cavity-amplified emission rate for inertial atoms can be further enhanced by centripetal acceleration.
When the atom is resonant with the cavity, i.e., when $\omega_{0}=\omega_{c}$, the emission rate for inertial atoms within the cavity reaches its maximum, given by the peak value:
@@
×$\times(d_{\rho}^{2}+d_{\phi}^{2}),$ (10)
where $Q\gg 1$. As the rotational angular velocity increases from zero, the contribution of the transverse polarization (atomic polarization perpendicular to the rotation axis) to the emission rate is strongly suppressed by rotation. Consequently, the primary contribution to the emission rate arises from the axial polarization (atomic polarization parallel to the rotation axis). According to Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), rotation always suppresses the emission rate, and the larger the rotational angular velocity, the smaller the emission rate becomes. This implies that, unlike in free space, the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by centripetal acceleration when the atom is resonant with the cavity.
When the atom is not resonant with the cavity, the emission rate, as described by Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), initially increases with the rotational angular velocity. After reaching a peak, it decreases as the rotational speed continues to increase. The peak appears at $\Omega=\omega_{c}-\omega_{0}$ when $\omega_{0}<\omega_{c}$, and at $\Omega=\omega_{0}-\omega_{c}$ when $\omega_{0}>\omega_{c}$. In either case, the peak value of the emission rate is
$\Gamma_{\downarrow}\approx\frac{Q}{6\epsilon_{0}\hbar V}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (11)
@@
Cavity effect.— Next, we investigate how the radiative properties of centripetally accelerated atoms can be significantly influenced by the presence of the cavity. To do this, we compare the peak values of the emission and excitation rates of these atoms when they are inside a cavity with those when they are in free space.
@@
1. We start with the emission rate. As shown in Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity), when the rotational angular velocity $\Omega$ exceeds the transition frequency of the atom $\omega_{0}$, i.e., when $\Omega>\omega_{0}$, the emission rate as a function of the cavity’s normal mode frequency exhibits two peaks, located at $\omega_{c}=\omega_{0}$ and $\omega_{c}=\omega_{0}+\Omega$, respectively. When $\Omega\leq\omega_{0}$, a third peak appears at $\omega_{c}=\omega_{0}-\Omega$. According to Eq. (Extensive manipulation of transition rates and substantial population inversion of rotating atoms inside a cavity) and Eq. (11), for isotropically polarizable atoms with $d_{z}^{2}=d_{\rho}^{2}=d_{\phi}^{2}\equiv d^{2}$, the leading terms of the peak emission rates at these points take the same value, which is $\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$. For comparison, the emission rate of a rotating atom in free space is yu24
$\Gamma_{\downarrow}^{\rm free}\approx$ $\frac{\omega_{0}^{3}}{3\pi\epsilon_{0}\hbar c^{3}}\left[\left(1+\frac{3\Omega^{2}}{\omega_{0}^{2}}\right)\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+d_{z}^{2}\right],$ (13)
when $\Omega\leq\omega_{0}$, and
$\Gamma_{\downarrow}^{\rm free}\approx$ $\frac{\left(\Omega+\omega_{0}\right){}^{3}}{6\pi\epsilon_{0}\hbar c^{3}}\left[\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\frac{2\omega_{0}^{3}}{\left(\Omega+\omega_{0}\right){}^{3}}d_{z}^{2}\right],$ (14)
when $\Omega>\omega_{0}$. Comparing $\Gamma_{\downarrow}=\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$ with Eq. (13) and Eq. (14) reveals that the emission rate of rotating atoms with isotropic polarizability can be enhanced by a factor of $\frac{Qc^{3}}{V\omega_{0}^{3}}$ when $\Omega\leq\omega_{0}$, and by a factor of $\frac{Qc^{3}}{V(\Omega+\omega_{0})^{3}}$ when $\Omega>\omega_{0}$, respectively. These enhancement factors can be extremely large, as $Q\gg 1$ and the cavity volume can be very small. As a result, the emission rate for rotating atoms inside the cavity is significantly amplified compared to that in free space yu24 .
@@
2. Next, we examine whether the excitation rate of centripetally accelerated atoms can also be significantly amplified by the presence of the cavity. We begin with the scenario when $\Omega>\omega_{0}$. According to Eq. (7), the excitation rate peaks when the normal mode frequency of the cavity is tuned to $\omega_{c}=\Omega-\omega_{0}$, with the peak value given by
$\Gamma_{\uparrow}\approx\frac{Q}{6\epsilon_{0}\hbar V}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (15)
which coincides with the peak value of the emission rate Eq. (11) at $\omega_{c}=\omega_{0}+\Omega$, indicating that the excitation rate for rotating atoms inside the cavity can be drastically amplified, even reaching the maximum cavity-amplified emission rate for rotating atoms. For comparison, the excitation rate of a centripetally accelerated atom in free space when $\Omega>\omega_{0}$ is given by yu24 ,
$\Gamma_{\uparrow}^{\rm free}\approx\frac{(\Omega-\omega_{0})^{3}}{6\pi\epsilon_{0}\hbar c^{3}}\left(d_{\rho}^{2}+d_{\phi}^{2}\right),$ (16)
Therefore, the maximum excitation rate of a centripetally accelerated atom inside a cavity is approximately $\frac{Qc^{3}}{V(\Omega-\omega_{0})^{3}}$ times that in free space.
On the other hand, when $\frac{\omega_{0}}{2}<\Omega\leq\omega_{0}$, the excitation rate of rotating atoms within a cavity reaches its peak when the normal mode frequency of the cavity is tuned to $\omega_{c}={2\Omega-\omega}_{0}$ according to Eq. (8), with the peak value given by the following expression:

### v2

@@ Abstract
We investigate the transition rates of a centripetally accelerated atom inside a high-quality cavity and show that they can be extensively tuned by adjusting the cavity resonance and the rotation frequency. Crucially, while inertial atoms cannot be excited in vacuum, rotation induces spontaneous excitation via the circular Unruh effect, with the cavity serving only as an amplifier. Using experimentally feasible parameters, we demonstrate that, in one scenario, the excitation rate can reach $\sim 10^{7}~\mathrm{s}^{-1}$ while emission remains negligible, enabling substantial population inversion. In another scenario, both excitation and emission can simultaneously attain $\sim 10^{7}~\mathrm{s}^{-1}$, corresponding to millions of transitions per second for a single atom. These findings highlight a powerful method for manipulating atomic transition rates for quantum applications and open a promising route toward experimental verification of the circular Unruh effect with state-of-the-art quantum technologies.
@@
Introduction.—The study of atomic transitions has long been central to our understanding of the interaction between matter and electromagnetic fields within the framework of quantum electrodynamics. In free space, an atom in its excited state can spontaneously emit a photon and transition to a lower energy state, a process known as spontaneous emission. Spontaneous emission in vacuum is governed by the interaction between the atom and the fluctuating electromagnetic field. Consequently, any modification of these vacuum fluctuations leads to changes in the atom’s radiative properties.
One particularly intriguing scenario arises when atoms are placed in a resonant cavity. The cavity modifies the density of states available for the fluctuating electromagnetic field, significantly enhancing the spontaneous emission rate in a phenomenon known as the Purcell effect Purcell46 , which has been experimentally observed Purcell2 ; Purcell3 and has played a pivotal role in advancing our understanding of how confined fluctuating electromagnetic fields influence atomic behavior.
@@
Vacuum fluctuations of the electromagnetic field can be profoundly altered not only by confinement to bounded spaces, such as cavities, but also by the noninertial motion of atoms. Atoms in such motion interact with these modified vacuum fluctuations in unique ways, leading to phenomena that have no counterpart in inertial motion. For instance, an atom undergoing uniform linear acceleration perceives the vacuum, as seen by an inertial observer, as a thermal bath. This leads to the atom spontaneously transitioning to a higher energy level, a phenomenon known as the Unruh effect Davies75 ; Unruh76 ; Fulling73 . Additionally, the emission rate of an accelerated atom is also enhanced compared to that of an inertial atom, mimicking the behavior of an atom in a thermal environment. However, the excitation rate typically remains much lower than the emission rate, except under extremely high accelerations where it can approach the emission rate. Therefore, while the Unruh effect is theoretically compelling, it has remained elusive in direct experimental observation due to the extreme accelerations required.
@@
Similar to the case of linear acceleration, an observer undergoing uniform centripetal acceleration would also perceive the vacuum of an inertial atom as an environment with radiation, although the spectrum of this radiation is nonthermal Bell83 ; Hacyan1986 ; Bell87 ; Kim1987 ; Unruh98 ; Rosu2005 . Consequently, spontaneous excitation can occur for atoms under centripetal acceleration in a vacuum, a phenomenon that can be regarded as the circular Unruh effect. The emission rate in this scenario can also be enhanced by centripetal acceleration, although the behavior differs significantly from that under uniform linear acceleration Letaw1981 ; Takagi1984 ; Letaw1980 ; Rogers88 ; Davies1996 ; Lorenci2000 ; YaoJin2014-1 ; YaoJin2014-2 ; Biermann2020 ; Bunney23 ; Bunney23-2 ; yu24 .
@@
Interestingly, recent studies have shown that when the rotational angular velocity exceeds the transition frequency of the atom, the excitation rate can reach the same order of magnitude as the emission rate, even when the orbital radius and thus the centripetal acceleration are extremely small yu24 . These findings not only challenge the conventional understanding that a significant Unruh effect requires a large acceleration and reveal that the radiative properties of a rotating atom can be profoundly different from those of inertial atoms, even at vanishingly small centripetal accelerations, but also naturally lead to further questions. Specifically, on one hand, how would the radiative properties of inertial atoms inside a cavity change if they are rotated? Can the maximum cavity-amplified emission rate for inertial atoms be further enhanced by centripetal acceleration, as it is in free space? On the other hand, how would the radiative properties of rotating atoms in free space change if they are placed inside a cavity? Can the excitation rate of noninertial atoms be drastically amplified by cavities, similar to how the emission rate is for inertial atoms? If possible, can the excitation rate even exceed the emission rate, leading to population inversion? These questions lie at the heart of this study, where we explore the interplay between the effects of atomic circular acceleration, a form of noninertial motion, and the confinement of the electromagnetic field within the cavity. Our goal is to control atomic transitions, which are crucial for developing high-brightness single-photon sources for quantum communication, improving the precision of atomic clocks for quantum metrology, advancing quantum computing through enhanced qubit readout speeds, and experimentally verifying the circular Unruh effect using state-of-the-art quantum technologies.
@@
Transition rates of centripetally accelerated atoms in a cavity.—We consider a two-level atom with a proper transition frequency $\omega_{0}$ coupled to fluctuating vacuum electromagnetic fields within a cavity. Assuming that the atom moves along a circular orbit, its trajectory in the laboratory frame is described by $x(t)=R\cos(\Omega t),~y(t)=R\sin(\Omega t),~z(t)=0$, where $R$ represents the radius of the orbit, $\Omega$ the rotational angular velocity, and $t$ the coordinate time. In the laboratory frame, the Hamiltonian characterizing the dipole interaction between the atom and the electromagnetic field can be expressed as $H_{I}=-\sum_{m=\rho,\phi,z}D_{m}\mathcal{E}_{m}$ yu24 , where $D_{m}$ is the electric dipole moment operator of the atom, and the components of $\mathcal{E}_{m}$ are
@@
Using perturbation theory, the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$ of the atom in the laboratory frame can be derived as follows:
$\Gamma_{\downarrow\uparrow}=\sum_{m,n=\rho,\phi,z}\frac{{d}_{m}{d}^{*}_{n}}{\hbar^{2}}\int dt_{-}e^{\pm i\omega t_{-}}G_{mn}(t_{-})\;,$ (2)
where $t_{-}=t-t^{\prime}$, $\hbar$ is the reduced Planck constant, ${d}_{m}=\langle e|D_{m}|g\rangle$ is the dipole transition matrix element with $|g\rangle$ ($|e\rangle$) being the ground (excited) state of the atom, $G_{mn}(t_{-})=\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{n}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$ are the two-point functions of the electromagnetic fields with $|0\rangle$ being the vacuum state of the electromagnetic fields, and $\omega=\omega_{0}(1-{\Omega^{2}R^{2}}/{c^{2}})^{\frac{1}{2}}$ is the transition frequency of the atom in the laboratory frame. The ++ and - signs in the exponential function correspond to the emission rate $\Gamma_{\downarrow}$ and excitation rate $\Gamma_{\uparrow}$, respectively. For the explicit forms of the two-point functions $\langle 0|\mathcal{E}_{m}\left(t,\mathbf{x}\right)\mathcal{E}_{m}\left(t^{\prime},\mathbf{x}^{\prime}\right)|0\rangle$, please refer to Sec. I of Supplemental Material. These functions are dependent on the explicit form of the density of states within the cavity $\rho(\omega_{k})$, where $\omega_{k}$ denotes the frequency of the field mode. The density of states within a cavity is typically modeled as a Lorentzian distribution given by $\rho(\omega_{k})=\frac{1}{\pi}\frac{\omega_{c}/Q}{(\omega_{c}/Q)^{2}+(\omega_{k}-\omega_{c})^{2}}$, where $Q~(Q\gg 1)$ represents the quality factor, and $\omega_{c}$ denotes the normal mode frequency of the cavity. By substituting the explicit forms of the field correlation functions into Eq. (2), the leading terms of the emission rate $\Gamma_{\downarrow}$ in the nonrelativistic limit ($v=R\Omega\ll c$) can be calculated as
@@
$\qquad+\left(d_{\rho}^{2}+d_{\phi}^{2}\right)\times\left[\frac{\omega_{0}+\Omega}{Q^{2}\left(\omega_{0}+\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right$.
$\qquad\left.+\frac{\left(\omega_{0}-\Omega\right)\Theta\left(\omega_{0}-\Omega\right)}{Q^{2}\left(\omega_{0}-\Omega-\omega_{c}\right){}^{2}+\omega_{c}^{2}}\right]\bigg\}+\mathcal{O}\left[({R\Omega}/{c}\right)^{2}].$~~~ (3)
@@
$\Gamma_{\uparrow}\approx\frac{Q\omega_{c}\left(\Omega-\omega_{0}\right)\left(d_{\rho}^{2}+d_{\phi}^{2}\right)}{6\epsilon_{0}\hbar V\left[Q^{2}\left(\Omega-\omega_{0}-\omega_{c}\right){}^{2}+\omega_{c}^{2}\right]}+\mathcal{O}\left[({R\Omega}/{c}\right)^{2}],$~ (4)
for $\Omega>\omega_{0}$. Here $\epsilon_{0}$ is the vacuum permittivity, $V$ represents the volume of the cavity, and $\Theta(x)$ is the Heaviside step function. Note that when $\Omega<\omega_{0}$, the leading term of the excitation rate is of the order $({R\Omega}/{c})^{2}$ or higher, as detailed in Sec. II of Supplemental Material. In the following, we discuss explicitly the properties of the transition rates for centripetally accelerated atoms within a cavity, focusing on the interplay between the effects of centripetal acceleration and cavity. It should be noted that the analysis assumes $Q\gg 1$.
Rotational effect—First, we examine the impact of rotation on the radiative properties of atoms by comparing the transition rate of centripetally accelerated atoms inside a cavity with that of inertial atoms in the same environment. Since inertial atoms in vacuum cannot be excited regardless of the cavity design, our focus is on the emission rate. In particular, we investigate whether the maximum cavity-amplified emission rate for inertial atoms can be further enhanced by centripetal acceleration.
Rotation-induced suppression at resonance: When the atom is resonant with the cavity, i.e., when $\omega_{0}=\omega_{c}$, the emission rate for inertial atoms within the cavity reaches its maximum, given by the peak value:
@@
$\times(d_{\rho}^{2}+d_{\phi}^{2})+\mathcal{O}[\left({R\Omega}/{c}\right)^{2}].$ (6)
As the rotational angular velocity increases from zero, the contribution of the transverse polarization (atomic polarization perpendicular to the rotation axis) to the emission rate is strongly suppressed by rotation. Consequently, the primary contribution to the emission rate arises from the axial polarization (atomic polarization parallel to the rotation axis). According to Eq. (Extensive Manipulation of Transition Rates and Substantial Population Inversion of Rotating Atoms Inside a Cavity), rotation always suppresses the emission rate, and the larger the rotational angular velocity, the smaller the emission rate becomes. This implies that, unlike in free space, the maximum cavity-amplified emission rate for inertial atoms cannot be further enhanced by centripetal acceleration when the atom is resonant with the cavity.
Rotation-induced enhancement off resonance: When the atom is not resonant with the cavity, the emission rate, as described by Eq. (Extensive Manipulation of Transition Rates and Substantial Population Inversion of Rotating Atoms Inside a Cavity), initially increases with the rotational angular velocity. After reaching a peak, it decreases as the rotational speed continues to increase. The peak appears at $\Omega=\omega_{c}-\omega_{0}$ when $\omega_{0}<\omega_{c}$, and at $\Omega=\omega_{0}-\omega_{c}$ when $\omega_{0}>\omega_{c}$. In either case, the peak value of the emission rate is
$\Gamma_{\downarrow}\approx\frac{Q}{6\epsilon_{0}\hbar V}\left(d_{\rho}^{2}+d_{\phi}^{2}\right)+\mathcal{O}[\left({R\Omega}/{c}\right)^{2}],$ (7)
@@
Cavity effect—Next, we investigate how the transitions of centripetally accelerated atoms can be significantly influenced by the presence of the cavity. To do this, we compare the peak values of the emission and excitation rates of these atoms when they are inside a cavity with those when they are in free space.
@@
Significant enhancement of emission rate: We start with the emission rate. As shown in Eq. (Extensive Manipulation of Transition Rates and Substantial Population Inversion of Rotating Atoms Inside a Cavity), when the rotational angular velocity $\Omega$ exceeds the transition frequency of the atom $\omega_{0}$, i.e., when $\Omega>\omega_{0}$, the emission rate as a function of the cavity’s normal mode frequency exhibits two peaks, located at $\omega_{c}=\omega_{0}$ and $\omega_{c}=\omega_{0}+\Omega$, respectively. When $\Omega\leq\omega_{0}$, a third peak appears at $\omega_{c}=\omega_{0}-\Omega$. According to Eq. (Extensive Manipulation of Transition Rates and Substantial Population Inversion of Rotating Atoms Inside a Cavity) and Eq. (7), for isotropically polarizable atoms with $d_{z}^{2}=d_{\rho}^{2}=d_{\phi}^{2}\equiv d^{2}$, the leading terms of the peak emission rates at these points take the same value, which is $\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$. Comparing $\Gamma_{\downarrow}=\frac{Qd^{2}}{3\epsilon_{0}\hbar V}$ with the emission rate of a rotating atom in free space as shown in Eqs. (15)-(17) of Ref. yu24 reveals that the emission rate of rotating atoms with isotropic polarizability can be enhanced by a factor of $\frac{Qc^{3}}{V\omega_{0}^{3}}$ when $\Omega\leq\omega_{0}$, and by a factor of $\frac{Qc^{3}}{V(\Omega+\omega_{0})^{3}}$ when $\Omega>\omega_{0}$, respectively. These enhancement factors can be extremely large, as $Q\gg 1$ and the cavity volume can be very small. As a result, the emission rate for rotating atoms inside the cavity is significantly amplified compared to that in free space yu24 .
@@
Significant enhancement of excitation rate: Next, we examine whether the excitation rate of centripetally accelerated atoms can also be significantly amplified by the presence of the cavity. According to Eq. (4), when $\Omega>\omega_{0}$, the excitation rate peaks when the normal mode frequency of the cavity is tuned to $\omega_{c}=\Omega-\omega_{0}$, with the peak value given by